Given the equation y=6x-2 what should the graph look like? I am not sure where to start in order to graph this to show what the graph should look like. Could someone please help?
Answer by checkley71(8403)

THE LINE REPRESENTED BY Y=6X-2 IS A LINE WITH A SLOPE OF 6 AND A Y INTERCEPT OF -2. SETTING Y=0 WE GET ANOTHER POINT 0=6X-2 OR 6X=2 OR X=2/6 OR X=1/3 THUS WE NOW HAVE TWO POINTS (0,-2)(1/3,0) AND A SLOPE OF 6.
PLOT THE TWO POINTS AND DRAW A LINE THROUGH THEM.